OpenStudy (anonymous):

What is the distance between E(5, 1) and F(7, 10)? Round to the nearest tenth. What is the distance between R(–5, 2) and S(4, –7)? Round to the nearest tenth. Can you give me the filled in formula again? @Michele_Laino

OpenStudy (michele_laino):

question #1 hint: the requested distance d is: \[\large d = \sqrt {{{\left( {5 - 7} \right)}^2} + {{\left( {1 - 10} \right)}^2}} = ...?\]

OpenStudy (anonymous):

The length of the slope is given by the square root of the difference in squares of the different of x and y values. or